OpenStudy (anonymous):

What is the value of the missing exponent in the problem a^2*a^5*a^3=a^? ?

OpenStudy (anonymous):

a^10

jimthompson5910 (jim_thompson5910):

When multiplying exponential expressions with the same base, you add the exponents. So the final exponent is 2+5+3 = 7+3 = 10

OpenStudy (anonymous):

Add the exponents so the answer is 10